OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 12x – y = –4 4x – 3y = –6 Answer 4x – 3(–12x – 4) = –6 4(12x + 4) – 3y = –6 4(–12x – 4) – 3y = –6 4x – 3(12x + 4) = –6

OpenStudy (anonymous):

12x – y = –4 rearranging this first one we get 12x + 4 = y substituting this in second equation we get 4x – 3(12x + 4) = –6 i.e the last one

OpenStudy (anonymous):

12x-y = -4 12x+4 = y 4x-3(12x+4)=-6 4x-36-12=-6 4x-48=-6 4x=42 x=14